# Direct Preference Optimization Needs

### v3

@@ B.1 Implementation Details
IPO (Azar et al., 2023) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
@@
$\alpha$-DPO $-\log\sigma\left(u(x,y_{w},y_{l})-\operatorname{sg}\left[\gamma+\alpha M^{\ast}(x,y_{w},y_{l})\right]\right)$ $\beta\in[2.5,10.0],\,\,\gamma\in[0.1,0.3,0.5]$
@@
Table 4: The hyperparameter values in $\alpha$-DPO used for each training setting.
@@
Hyperparameter in $\alpha$-DPO.
Table 4 outlines the hyperparameters used for $\alpha$-DPO under various settings. It’s important to note that while our approach involves three key parameters, we have found through experience that $\beta$ can be reliably set to 10.0 by default. Among these parameters, $\gamma$ typically requires more careful tuning. As for $\alpha$, we have observed consistent performance improvements when set to 5e-2 by default. If you are already familiar with the parameter settings for SimPO, you can focus your search primarily on $\alpha$ or simply adopt the default setting of $\alpha=5e-2$.
@@
B.2 $\alpha$-DPO without length-normalized
In this paper, we consider length-normalized training as a stability technique and not as a primary contribution of this work. Existing research (Meng et al., 2024) has demonstrated that length normalization can indeed enhance model performance, particularly with respect to the length control win rate. However, to validate the general applicability of $\alpha$-DPO—specifically, its stability and performance without length normalization—we conducted experiments across several models: meta-llama/Meta-Llama-3-8B-Instruct, mistralai/Mistral-7B-Instruct-v0.2, and google/gemma-2-9b-it.
Figure 4: $\alpha$-DPO LC on AlpacaEval 2 with different $\alpha$ values.
We evaluated DPO, SimPO without length normalization, and $\alpha$-DPO without length normalization. The experimental results, as shown in Table 5, demonstrate that $\alpha$-DPO consistently achieves performance improvements even without the use of length normalization. This indicates the robustness and general effectiveness of $\alpha$-DPO.
@@ B.2 $\alpha$-DPO without length-normalized
$\alpha$-DPO w/o LN 44.4 42.6 32.0 38.4 51.1 54.0 72.7 70.5
B.3 $\alpha$-DPO with differenct $\alpha$
@@
It is noteworthy that within the parameter tuning range [1e-2, 5e-2, 0.1, 0.2], the optimal $\alpha$ values are consistently around 0.1 or even closer to 5e-2. This observation aligns with our Lemma 3.2, which indicates that smaller $\alpha$ values result in a lower estimation error in the online SimPO.
B.4 Comparison With TDPO
To investigate the relationship between TDPO and $\alpha$-DPO, we conducted the experiments, with the results outlined below.
Table 6: Performance comparison between TDPO and $\alpha$-DPO.
@@ B.4 Comparison With TDPO
$\alpha$-DPO w/ $\delta(x,y_{w},y_{l})$ 56.9 50.4
$\alpha$-DPO w/ ${M}(x,y_{w},y_{l})$ 58.7 51.1
In its original form, TDPO did not perform well on LLAMA2-8B. By applying Lemma 3.3, we modified the expression $M(x,y_{w},y_{l})$ in $\alpha$-DPO to use TDPO’s $\delta(x,y_{w},y_{l})$, converting our sentence-level estimations to a token-level calculation. This adjustment resulted in a noticeable performance improvement, which we attribute to the length-normalization, $\gamma$ and z-score normalization of $\delta(x,y_{w},y_{l})$. Nevertheless, the modified TDPO still underperformed compared to $\alpha$-DPO. This indicates that, when the $\pi_{\text{ref}}$ is suboptimal, token-level calculations are prone to significant errors.
Figure 5: (a) SimPO: Reward difference distribution under different $\gamma$ values. (b) $\alpha$-DPO: Reward difference distribution under different $\alpha$ values. (c) SimPO: Log likelihood distribution on chosen responses under different $\gamma$ values. (d) $\alpha$-DPO: Log likelihood distribution on chosen responses under different $\alpha$ values.

### v4

@@ D.1 Implementation Details
IPO (Azar et al., 2024) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
@@
AlphaDPO $-\log\sigma\left(u(x,y_{w},y_{l})-\operatorname{sg}\left[\gamma+\alpha M^{\ast}(x,y_{w},y_{l})\right]\right)$ $\beta\in[2.5,10.0],\,\,\gamma\in[0.1,0.3,0.5]$
@@
Table 4: The hyperparameter values in AlphaDPO used for each training setting.
@@
Hyperparameter in AlphaDPO.
Table 4 outlines the hyperparameters used for AlphaDPO under various settings. It’s important to note that while our approach involves three key parameters, we have found through experience that $\beta$ can be reliably set to 10.0 by default. Among these parameters, $\gamma$ typically requires more careful tuning. As for $\alpha$, we have observed consistent performance improvements when set to 5e-2 by default. If you are already familiar with the parameter settings for SimPO, you can focus your search primarily on $\alpha$ or simply adopt the default setting of $\alpha=5e-2$.
@@
D.2 AlphaDPO without length-normalized
In this paper, we consider length-normalized training as a stability technique and not as a primary contribution of this work. Existing research (Meng et al., 2024) has demonstrated that length normalization can indeed enhance model performance, particularly with respect to the length control win rate. However, to validate the general applicability of AlphaDPO—specifically, its stability and performance without length normalization—we conducted experiments across several models: meta-llama/Meta-Llama-3-8B-Instruct, mistralai/Mistral-7B-Instruct-v0.2, and google/gemma-2-9b-it.
Figure 4: AlphaDPO LC on AlpacaEval 2 with different $\alpha$ values.
We evaluated DPO, SimPO without length normalization, and AlphaDPO without length normalization. The experimental results, as shown in Table 5, demonstrate that AlphaDPO consistently achieves performance improvements even without the use of length normalization. This indicates the robustness and general effectiveness of AlphaDPO.
@@ D.2 AlphaDPO without length-normalized
AlphaDPO w/o LN 44.4 42.6 32.0 38.4 51.1 54.0 72.7 70.5
D.3 AlphaDPO with differenct $\alpha$
@@
It is noteworthy that within the parameter tuning range [1e-2, 5e-2, 0.1, 0.2], the optimal $\alpha$ values are consistently around 0.1 or even closer to 5e-2.
D.4 Comparison With TDPO
To investigate the relationship between TDPO and AlphaDPO, we conducted the experiments, with the results outlined below.
Table 6: Performance comparison between TDPO and AlphaDPO.
@@ D.4 Comparison With TDPO
AlphaDPO w/ $\delta(x,y_{w},y_{l})$ 56.9 50.4
AlphaDPO w/ ${M}(x,y_{w},y_{l})$ 58.7 51.1
In its original form, TDPO did not perform well on Llama3-8B. By applying Lemma 3.1, we modified the expression $M(x,y_{w},y_{l})$ in AlphaDPO to use TDPO’s $\delta(x,y_{w},y_{l})$, converting our sentence-level estimations to a token-level calculation. This adjustment resulted in a noticeable performance improvement, which we attribute to the length-normalization, $\gamma$ and z-score normalization of $\delta(x,y_{w},y_{l})$. Nevertheless, the modified TDPO still underperformed compared to AlphaDPO. This indicates that, when the $\pi_{\text{ref}}$ is suboptimal, token-level calculations are prone to significant errors.
Figure 5: (a) SimPO: Reward difference distribution under different $\gamma$ values. (b) AlphaDPO: Reward difference distribution under different $\alpha$ values. (c) SimPO: Log likelihood distribution on chosen responses under different $\gamma$ values. (d) AlphaDPO: Log likelihood distribution on chosen responses under different $\alpha$ values.